Please help me I don't want to flunk.1. The function H described by H(x) =2.75x+71.48 can be used to predict the height, in centimeters, of a woman whose humorous (the bone from the elbow to the shoulder) is x cm long.
Predict the height of a woman whose numerous is 33cm long.
The height of a woman whose numerous is 33 cm long is __cm.
2. Train A and B are traveling in the same direction on parallel tracks. Train A is traveling at 60 miles per hour and B is traveling at 80 miles per hour. Train A passes a station at 12:20 A.M. If train B pass the same station at 12:35 A.M. at what time will train B catch up with train A?
When will B catch up with Train A? __:__ __

3. In 1933, the life expectancy of males in a certain country was 62.3 years. In 1998, it was 65.9 years let E represent the life expectancy in year t and let t represent the number of years since 1993.
The linear function E (t) that fits the data is E (t) =__t +__
Round to the nearest tenth
Use the function to predict the life expectancy of males in 2005 E (12) =
(Round to the nearest tenth)
4. 9x-9y=63
9y-9x=-63 Could I say that there is no solution for this system.

Answer by stanbon(75887)

1. The function H described by H(x) =2.75x+71.48 can be used to predict the height, in centimeters, of a woman whose humorous (the bone from the elbow to the shoulder) is x cm long.
Predict the height of a woman whose numerous is 33cm long.
The height of a woman whose numerous is 33 cm long is __cm.
H(33) = 2.75*33 + 71.48 = 162.23 cm
------------------------------------------------
2. Train A and B are traveling in the same direction on parallel tracks. Train A is traveling at 60 miles per hour and B is traveling at 80 miles per hour. Train A passes a station at 12:20 A.M. If train B pass the same station at 12:35 A.M. at what time will train B catch up with train A?
When will B catch up with Train A? __:__ __
Train A DATA:
rate = 60 mph ; time = x hrs ; distance = rt = 60x miles
---------------
Train B DATA:
rate = 80 mph ; time = (x-15/60) hrs ; distance = 80(x-(1/4)) miles
--------------------
Equation:
distance A = distance B
60x = 80(x-(1/4))
60x = 80x - 20
20x = 20
x = 1 hr
B catches A at 1:20 AM
------------------------------------
3. In 1993, the life expectancy of males in a certain country was 62.3 years. In 1998, it was 65.9 years let E represent the life expectancy in year t and let t represent the number of years since 1993.
The linear function E (t) that fits the data is E (t) =__t +__
Round to the nearest tenth
Use the function to predict the life expectancy of males in 2005 E (12) =
(Round to the nearest tenth)
---
Let t=0 correspond to 1993; then t=5 corresponds to 1998.
You have two points: (0,62.3) and (5,65.9)
slope = (65.9-62.3)/(65-0) = 0.72
intercept:62.3
Equation: E(t) = 0.72t + 62.3
E(12) = 0.72*12 + 62.3 = 70.94
---------------
4. 9x-9y=63
9y-9x=-63 Could I say that there is no solution for this system.
Rearrange to get:
9x - 9y = 63
-9x+ 9y = -63
-----------------
Add to get
0 = 0
This means the system is true for any value of x and that y = x-7
The two equations are really just one equation: y = x - 7
It has an infinite number of solutions.
=============================================================
Cheers,
Stan H.